OpenStudy (anonymous):

3 times the complement of an angle is 30 more than the supplement of the original angle. What is the measure of the angle?

OpenStudy (doc.brown):

What's a complement of an angle?

OpenStudy (anonymous):

= 90 degrees

OpenStudy (doc.brown):

Supplement?

OpenStudy (anonymous):

180

OpenStudy (doc.brown):

\[3(90)+30=180x\]

OpenStudy (doc.brown):

or is the supplement added?

OpenStudy (anonymous):

The complement of an angle is (90-x), not just 90, right?

OpenStudy (anonymous):

I'm pretty sure of this.

OpenStudy (anonymous):

Also, supplement of an angle is (180-x) not just 180, where x is the angle.

OpenStudy (doc.brown):

ah, that makes sense. listen to @mathbrz

OpenStudy (anonymous):

\[3(90-x) = (180-x) + 30\]

OpenStudy (anonymous):

so 60 degrees?

OpenStudy (anonymous):

I got x = 30.

OpenStudy (doc.brown):

This part I do know,\[3(90-x)=(180-x)+30\]\[3(90)+3(-x)=180-x+30\]\[270-3x=210-x\]\[270-210=3x-x\]\[2x=60\]\[x=30\]

OpenStudy (anonymous):

ok thanks i see what i did wrong

OpenStudy (anonymous):

gj

OpenStudy (anonymous):

Thanks for your help @doc.brown @mathbrz

OpenStudy (anonymous):

yw :) have a nice evening. going to bed